# Rethinking Visual Dependency in Long-Context Reasoning for Large Vision-Language Models

### v1

@@ 6.1 Experimental Setting
To validate the effectiveness of our proposed method, we conduct experiments using the long-context dataset collected as mentioned in Sec. 4.1, which is based on the SVIT dataset (Zhao et al., 2023). It comprises 15,295 samples, divided into five groups with 0.5k, 1k, 1.5k, 2k, and 2.5k samples, respectively. The LVLMs with 7B parameters used in our experiments include DeepSeek-VL, LLaVA, LLaVA-Next, ViP-LLaVA, LLaVA-Next-Video, and Video-LLaMA2. The token pruning is applied to the first 10 layers, with a pruning rate of 0.3. For evaluation, we utilize greedy decoding for LVLMs and conduct model inference on NVIDIA A100 GPUs.
@@ 6.2 Results
Figure 7: Top: Performance comparison of our method and various LVLMs. Bottom: Performance gain of our method compared with baselines.
In Figure 7, the performance of different LVLMs is compared across varying context lengths, with the top figure showing the performance of various LVLMs with our method and the bottom figure showing the performance gain of our method compared with baselines. Our method demonstrates strong and consistent performance, especially with Video-LaMA2, which maintains high performance across contexts of different lengths, indicating robustness to different context lengths. Although most models show a slight drop in performance as the context length increases, the performance gain analysis reveals that our method is especially effective in improving models that struggle with handling long contexts, with more significant gains observed for models with lower baseline performance. Even if models that already perform well at baseline show improvements, suggesting our method can provide refinement for LVLMs under long-context reasoning. In addition, due to token pruning, our method also offers a significant advantage in inference efficiency, reducing inference time to approximately half of the original.
@@ 6.3 Analysis on Token Pruning
Figure 8: Results with different token pruning methods.
@@
Figure 8 shows the performance comparison of different token pruning methods: “Max Pooling”, “Mean Pooling”, and “Random”. Both “Max Pooling” and “Mean Pooling” maintain high accuracy as the sequence length decreases, supporting our claim in Proposition 1 that pruning low-importance tokens is crucial. These methods leverage multi-head attention to preserve key tokens with higher attention weights, ensuring minimal performance loss. In contrast, “Random” leads to a complete loss in accuracy, highlighting the importance of intelligently selecting tokens for pruning to maintain model performance.
Differences in Token Pruning Layers.
@@
Impact of Token Pruning layer Numbers.
Figure 10: Results with different token pruning layer numbers.
@@
Impact of Token Pruning Rate.
@@
Pruned Token Distribution.
@@
6.4 Preliminary Exploration on Token Pruning Scaling Laws
@@
6.5 Case Study
7 Conclusion
In this study, we investigate the challenges faced by LVLMs in long-context reasoning. Our empirical analysis demonstrated that LVLMs experience a significant decline in visual dependency as the context length increases. We also identified that a higher proportion of target objects within images helps stabilize visual dependency and that cross-modal interactions are more prominent in the shallow layers of LVLMs, shifting towards predominantly textual interactions in deeper layers. To mitigate the performance degradation, we proposed a novel context pruning method that strategically removes less critical textual information. Our approach was validated using a newly constructed long-context dataset based on SVIT, and it proved effective across various LVLMs. Additional experiments confirmed the robustness of different token pruning strategies and elucidated the relationship between pruning rates and context length.
Limitations
While this study provides significant insights into the performance of LVLMs in long-context reasoning, there is still room for further research. The extensive computational resources required for our analysis limited the number of LVLMs we were able to evaluate. Additionally, our investigation into Token Pruning Scaling Laws was conducted as a preliminary exploration due to resource constraints. Although we identified a relationship between context length and Token Pruning rate, more comprehensive studies are needed to fully elucidate this connection and to establish robust scaling laws.

### v2

@@ 5.1 Experimental Setting
To validate the effectiveness of our proposed method, we conduct experiments using the long-context dataset collected as mentioned in Sec. 3.1, which is based on the SVIT dataset (Zhao et al., 2023). It comprises 15,295 samples, divided into five groups with 0.5k, 1k, 1.5k, 2k, and 2.5k samples, respectively. The LVLMs with 7B parameters used in our experiments include DeepSeek-VL, LLaVA, LLaVA-Next, ViP-LLaVA, LLaVA-Next-Video, and Video-LLaMA2. The token pruning is applied to the first 10 layers, with a pruning rate of 0.3. For evaluation, we utilize greedy decoding and conduct model inference on NVIDIA A100 GPUs.
@@ 5.2 Results
In Table 1, we present the performance of different LVLMs across various context lengths, comparing the baseline methods with our approach. Our method demonstrates consistent improvements across all models, particularly for those struggling with long-context reasoning, such as DeepSeek-VL and ViP-LLaVA, where significant performance gains are observed. However, for models like LLaVA-Next and LLaVA-Next-Video, which already perform well at baseline, the improvements are relatively minor. To further analyze the effectiveness of our approach in such scenarios, we conducted additional experiments on the MMDU benchmark, as shown in Table 2. These results reveal that our method provides consistent yet modest gains, especially for highly optimized models, highlighting its capability to refine performance even when baseline results are already strong. Additionally, due to token pruning, our method significantly improves inference efficiency, reducing inference time by approximately half compared to the original setup. Moreover, Figure 8 shows that as the LLaVA-Next model size increases, the performance gap between our method and the baseline widens, demonstrating the scalability and effectiveness of our approach on larger models. In Table 3, our method achieves the highest average performance (44.3) on the MMDU benchmark with LLaVA-Next, outperforming both SpAtten Wang et al. (2021) (43.5) and LTP Bonnaerens and Dambre (2023) (44.0). This demonstrates our method’s superior ability to refine performance and handle contextual information more effectively than these alternatives. Details of MMDU dataset and more results can be found in Appendix D.
@@ 5.3 Analysis on Token Pruning
Figure 9: Comparison of results obtained using various token pruning methods. The random pruning approach ensures the final question token remains intact.
@@
Figure 9 shows the performance comparison of different token pruning methods: “Max Pooling”, “Mean Pooling”, and “Random”. Both “Max Pooling” and “Mean Pooling” maintain high accuracy as the sequence length decreases, supporting our claim in Proposition 1 that pruning low-importance tokens is crucial. These methods leverage multi-head attention to preserve key tokens with higher attention weights, ensuring minimal performance loss. In contrast, “Random” leads to a complete loss in accuracy, highlighting the importance of intelligently selecting tokens for pruning to maintain model performance. More analysis on “Differences in Token Pruning Layers”, “Impact of Token Pruning layer Numbers”, and “Impact of Token Pruning Rate” can be found in Appendix E, Appendix F, and Appendix G, respectively.
Pruned Token Distribution.
Figure 10 illustrates the distribution of token types retained after pruning across model layers. While the overall proportion of token types remains fairly uniform, parts of speech show differing retention patterns. In contrast, Figure 9 highlights the severe performance drop caused by random pruning, which, despite similar pruning ratios, fails to retain critical information. Our method, as evidenced in Figure 10, effectively identifies and removes less important tokens, preserving key ones and maintaining model performance even with significant pruning. The detailed analysis can be found in the Appendix H. More analysis on “Token Pruning Scaling Laws”, “Pruned Token Position”, and “Case Study” can be found in Appendix I, Appendix J, and Appendix K, respectively.
6 Conclusion
In this study, we investigate the challenges faced by LVLMs in long-context reasoning. Our empirical analysis demonstrated that LVLMs experience a significant decline in visual dependency as the context length increases. We also identified that cross-modal interactions are more prominent in the shallow layers of LVLMs, shifting towards predominantly textual interactions in deeper layers. To mitigate the performance degradation, we proposed a novel context pruning method that strategically removes less critical textual information. Our approach was rigorously validated through extensive experiments, demonstrating its effectiveness across various LVLMs.
Limitations
While this study provides significant insights into the performance of LVLMs in long-context reasoning, there is still room for further research. The extensive computational resources required for our analysis limited the number of LVLMs we were able to evaluate. In addition, our investigation into Token Pruning Scaling Laws was conducted as a preliminary exploration due to resource constraints. Although we identified a relationship between context length and Token Pruning rate, more comprehensive studies are needed to elucidate this connection fully and to establish robust scaling laws.
Appendix A Related Work
A.1 Large Vision-Language Models
Recent advancements in LVLMs have focused on enhancing cross-modal understanding by integrating visual and textual inputs for complex tasks (Wang et al., 2023b; Chen et al., 2023a, b; Qu et al., 2024). For instance, LLaVA (Liu et al., 2023) leverages GPT-4 (OpenAI, 2023) to generate multimodal instruction-following data, improving performance in tasks like Science QA and multimodal reasoning. Expanding on this, models such as SVIT (Zhao et al., 2023) and LLaVA-NeXT (Liu et al., 2024a) utilize larger datasets and improve capabilities in OCR and abstract reasoning. Approaches like DeepSeek-VL (Lu et al., 2024) address real-world challenges by handling noisy and ambiguous inputs, while ViP-LLaVA (Cai et al., 2024) enhances robustness in processing diverse visual prompts. Additionally, some methods such as SURf’s self-refinement framework improve the accuracy of information retrieval (Sun et al., 2024), and VICL (Zhou et al., 2024) enables few-shot learning through inference on visual captioning. Zhang et al. (2024) extends to zero-shot video understanding using AnyRes, allowing models trained on images to generalize to videos, while Video-LLaMA2  (Cheng et al., 2024) enhances spatial-temporal modeling by integrating visual and audio inputs. These developments highlight LVLMs’ trends in scaling data, improving reasoning, and increasing robustness.
A.2 Long-Context Reasoning
Handling long contexts is a key challenge for LLM. Traditional position encoding techniques often struggle with long-range dependencies or varying-length sequences. To address this, methods like Contextual Position Encoding (CoPE) (Golovneva et al., 2024) dynamically adjust positional encodings based on context, improving performance on tasks requiring long-range dependencies. Training-free techniques, such as contextual sliding windows (An et al., 2024), extend the context length of pre-trained models without retraining, reducing computational overhead while maintaining sequence-processing capabilities. Parallel Context Windows (PCW) (Ratner et al., 2023) further enhances efficiency by splitting sequences into chunks for parallel processing. However, models tend to underutilize information from earlier inputs, favoring the middle or end (Liu et al., 2024b). Some methods like ThoT (Zhou et al., 2023) address this by improving reasoning across fragmented contexts, while LLMLingua (Jiang et al., 2023) compresses prompt representations, striking a balance between efficiency and context retention. Similarly, SpAtten (Wang et al., 2021) proposes an efficient sparse attention mechanism that leverages cascade token and head pruning to dynamically reduce computational overhead, significantly improving the model’s performance. LTMP (Bonnaerens and Dambre, 2023) introduces learned thresholds for token merging and pruning in Vision Transformers, achieving high token reduction rates with minimal accuracy loss by learning optimal trade-offs between merging and pruning across layers. These approaches showcase promising strategies to manage computational complexity while maintaining model performance across diverse tasks.
Appendix B Detail Analysis of Consistency Between Vision and Language Priors
To investigate the prediction consistency between vision and language priors, we analyze the disjoint portions of predictions with and without visual priors. Figure 3 presents the average performance of six models on different context lengths, comparing the accuracy of models provided visual priors (VL correct) and those only using language priors (Text correct). The proportion of samples that were correctly answered only when provided with an image generally decreases as the context length increases. It indicates that while visual dependency is more pronounced in shorter contexts, its impact diminishes as the model is exposed to more textual information. This suggests that the model increasingly relies on language reasoning under longer contexts. The proportion of samples only correctly answered without an image remains near zero across all context lengths. The results show that samples correctly answered using only language information remained to be correct even after adding visual information. It suggests that the models have effectively leveraged language priors to make accurate predictions, regardless of the availability of visual cues. As the context length increases, models rely more on the rich semantic information from text, reducing visual dependency.
Model Method C R VP LC AA IRU Avg.
LLaVA-Next-Video Zhang et al. (2024) Baseline 38.5 39.4 42.1 57.6 45.8 40.7 44.0
Ours 40.9 42.0 42.8 58.0 49.5 43.5 46.1
Video-LLaMA2 Cheng et al. (2024) Baseline 37.7 39.3 41.6 57.2 45.9 40.2 43.7
Ours 39.7 39.4 43.3 58.4 49.6 42.8 45.5
LLaVA-Next Liu et al. (2024a) Baseline 38.2 39.3 41.7 57.6 45.8 40.5 43.9
Ours 38.8 39.8 42.0 57.9 46.1 41.1 44.3
Table 4: Performance comparison between baselines and our method on MMDU.
Size Method C R VP LC AA IRU Avg.
7B Baseline 38.2 39.3 41.7 57.6 45.8 40.5 43.9
Ours 38.8 39.8 42.0 57.9 46.1 41.1 44.3
13B Baseline 38.4 39.7 41.8 58.0 46.5 40.8 44.2
Ours 40.8 41.2 43.6 58.3 46.9 42.8 45.6
34B Baseline 38.7 40.1 42.5 58.3 46.9 40.8 44.6
Ours 40.7 43.3 45.9 60.4 50.3 41.2 47.0
Table 5: Performance comparison across different model sizes (7B, 13B, 34B) of LLaVA-Next on MMDU.
Method C R VP LC AA IRU Avg.
Baseline 38.2 39.3 41.7 57.6 45.8 40.5 43.9
SpAtten Wang et al. (2021) 38.0 38.5 41.5 57.2 45.7 40.4 43.5
LTP Bonnaerens and Dambre (2023) 38.3 38.9 42.1 57.7 46.0 40.8 44.0
Ours 38.8 39.8 42.0 57.9 46.1 41.1 44.3
Table 6: Performance comparison of our method with SpAtten and LTP on MMDU, using the LLaVA-Next.
Appendix C Further Analysis of Information Flow
Figure 11 and Figure 7
Figure 11: Information flow results from baseline.
In this section, we extend the analysis of information flow by comparing the baseline results (Figure 11) with the pruned model (Figure 7). From Figure 11, it is evident that the baseline model distributes information flow more evenly across preserved tokens ($S_{vr}$), pruned tokens ($S_{vc}$), and the target ($S_{vt}$). However, the significance scores of $S_{vr}$ and $S_{vt}$ remain relatively lower in comparison to the pruned model. This indicates that, in the absence of pruning, the model’s attention is less focused, leading to higher entropy and a less stable flow of information. In contrast, the pruned results in Figure 7 show a noticeable increase in $S_{vr}$ and $S_{vt}$, alongside a significant drop in $S_{vc}$. This shift reflects the model’s improved ability to concentrate attention on essential tokens, aligning with Proposition 1, which states that pruning reduces entropy by stabilizing and concentrating attention. Moreover, the consistent rise in $S_{vr}$ and $S_{vt}$ across deeper layers highlights the model’s increasing reliance on visual inputs for decision-making, supporting Proposition 2.
Appendix D More Results on MMDU
This section presents detailed experimental results on the MMDU dataset, which is designed to evaluate multimodal capabilities across six metrics: Creativity (C), Richness (R), Visual Perception (VP), Logical Coherence (LC), Answer Accuracy (AA), and Image Relationship Understanding (IRU). Each of these metrics provides a comprehensive assessment of a model’s ability to generate coherent and meaningful responses in tasks involving multimodal data. The MMDU dataset consists of multimodal tasks where each sample we used includes two images, testing a model’s ability to process and relate visual inputs effectively. Performance evaluation follows the methodology outlined by Liu et al. (2024c), utilizing a set of standardized evaluation prompts. All models were assessed using GPT-4o, which serves as the scoring framework to ensure consistency and robustness in the evaluation process. Tables 4, 5, and 6 summarize the results of our method compared to baselines and prior approaches. Key findings are highlighted below:
Comparison with Baselines.
As shown in Table 4, our approach consistently outperforms baseline methods across all six metrics for different models, including LLaVA-Next-Video Zhang et al. (2024), Video-LLaMA2 Cheng et al. (2024), and LLaVA-Next Liu et al. (2024a). Notably, the average performance improvement ranges from 0.4 to 2.1 points, demonstrating the efficacy of our proposed enhancements.
Effect of Model Size.
Table 5 presents a detailed breakdown of performance for different model sizes (7B, 13B, and 34B). Larger models, such as the 34B version, show significant improvements, particularly in Visual Perception (VP) and Logical Coherence (LC), where our method achieves up to 2.4 points of improvement over the baseline. These results underscore the scalability of our approach as model size increases.
Comparison with Prior Methods.
In Table 6, we compare our method with SpAtten Wang et al. (2021) and LTP Bonnaerens and Dambre (2023) using the LLaVA-Next model. Our method achieves the highest scores across all metrics, with an average performance gain of 0.3 points compared to LTP, which is the closest competitor.
Appendix E Differences in Token Pruning Layers.
@@ Appendix D More Results on MMDU
Appendix F Impact of Token Pruning layer Numbers.
Figure 13: Performance of token pruning with different layer numbers.
@@
Appendix G Impact of Token Pruning Rate.
@@
Appendix H Detailed Analysis on Pruned Token Distribution
@@
Appendix I Token Pruning Scaling Laws
@@
Appendix J Pruned Token Position
Figure 17: Token retain rates across different length intervals for pruned tokens using our method.
Figure 17 shows the retained rate of tokens across different length intervals within the context. Our method demonstrates relatively uniform pruning across the earlier intervals (0–8), while tokens in the final interval (9) are retained at a significantly higher rate. Because the final interval often contains the question tokens, which are essential for generating accurate responses. The even pruning of intervals ensures efficient token usage, preserving essential information to improve model performance.
Appendix K Case Study